For what value of k does kx + 7 = 10 have a slope of 3?
Found 2 solutions by rapaljer, dabanfield:
Answer by rapaljer(4671)

Is there supposed to be a y somewhere in this equation??

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida

Answer by dabanfield(803)

The slope intercept form of the equation of a line is:
Y = M*X + B where M is the slope of the line and B is the Y intercept (the value of Y when the line crosses the X-axis).
In this case we have a line of the form 10 = K*X + 7 which is already in slope intercept form with K being the slope of the line. So, for the line to have a slope of 3, K must have the value 3.
The equation of the line with slope 3 then becomes 10 = 3*X + 7.